# Risk-Aware Autonomous Driving for Linear Temporal Logic Specifications

### v1

@@ Abstract
Decision-making for autonomous driving incorporating different types of risks is a challenging topic. This paper proposes a novel risk metric to facilitate the driving task specified by linear temporal logic (LTL) by balancing the risk brought up by different uncertain events. Such a balance is achieved by discounting the costs of these uncertain events according to their timing and severity, thereby reflecting a human-like awareness of risk. We have established a connection between this risk metric and the occupation measure, a fundamental concept in stochastic reachability problems, such that a risk-aware control synthesis problem under LTL specifications is formulated for autonomous vehicles using occupation measures. As a result, the synthesized policy achieves balanced decisions across different types of risks with associated costs, showcasing advantageous versatility and generalizability. The effectiveness and scalability of the proposed approach are validated by three typical traffic scenarios in Carla simulator.
@@ I Introduction
Developing motion-planning frameworks for autonomous vehicles remains a significant challenge due to the inherent complexity and uncertainty of real-world traffic scenarios [1, 2]. The complexity primarily arises from the diverse and comprehensive concerns in a practical driving task, including safety [3], traffic rules [4], and social norms [5, 6]. The uncertainty mainly originates from the dynamic environment with unpredictable traffic participants [7] and unexpected events [8]. Taking an unprotected turn scenario [9] in Fig. 1 as an example, the ego vehicle must take a left turn to reach a predefined goal without collisions with the opponent car. It must obey the traffic rules by correctly reacting to the traffic light, rendering a motion planning problem subject to hard spatiotemporal constraints [10]. Moreover, it should also follow certain driving norms, such as giving priority to an opponent car driving from the opposite direction. The environmental uncertainty brings additional challenges to the driving task by imposing the ego vehicle to the potential risk of violating the specifications above, even rendering the task infeasible in extreme cases [11]. A risk-aware autonomous driving scheme requires the ego vehicle to make decisions incorporating different risks associated with safety, traffic rules, and social norms, as mentioned above.
@@
Temporal logic provides a formal and expressive framework to describe high-level tasks that incorporates both spatial and temporal constraints [12, 13] and synthesize provably correct controllers [14], making itself well-suited to encode traffic tasks with rules and norms [15] and enhance trustworthiness in motion planning tasks [16, 17]. For motion planning in an uncertain environment, risk can be quantified as the probability that a temporal logic specification fails [18]. To develop a risk-aware control policy, many control synthesis problems focus on maximizing the satisfaction probability of temporal logic specifications, i.e., minimizing the violation risk [19, 20, 21]. However, such risk metrics are not suitable for realistic traffic where different types of risky events may be involved [22]. In this sense, a rational policy should prioritize events according to their emergency and timing [23]. For example, the dangers in the near future should be prioritized compared to those in the far future [24]. Unfortunately, a single satisfaction probability of temporal logic specifications in most previous methods cannot differentiate the severity and timing of events [25], potentially resulting in unsafe driving behavior. Thus, a novel risk-aware motion planning method is needed to balance different types of risk according to the severity and timing of various events [26], ensuring more reliable and human-like decision-making in traffic scenarios [27].
A model commonly used in the transportation field, named driver’s risk field [24, 28], has been designed to mimic a human-like awareness of risk by incorporating discounting factors and cost functions for different events [29]. We believe similar techniques can also be used to address the above limitations of satisfaction probability of temporal logic. Notably, we found the connection between the risk field model and the occupation measure, a key concept for solving a stochastic reachability problem for a Markov Decision Process (MDP) [30]. The occupation measure has been used as heuristics to guide an efficient search for motion planning problems [31]. In [32], it is used to synthesize policies that balance different probabilities of temporal logic specifications for Markov decision process (MDP) models by a coupled linear programming (LP) problem, providing an efficient solution for control synthesis with balanced risk. In [33], it has been used to derive a robust control policy with probabilistic guarantees, addressing a correct-by-design solution for systems with stochastic uncertainty. Inspired by this, we propose to incorporate the risk field model into the control synthesis problem using occupation measures.
This paper proposes a novel risk-aware motion planning framework for autonomous vehicles incorporating different types of risk in a dynamic and uncertain environment. Firstly, we employ a combination of safety and co-safety specifications, which are LTL fragments [34], to represent traffic tasks without losing expressivity. Such specifications allows to convert the control synthesis problem into a reachability problem, which avoid expensive computations. Secondly, we develop a novel risk metric for LTL specifications, which accounts for the timing and severity of different uncertain events. In addition, we establish the connection between this risk metric and the occupation measure dedicated to a reachability problem. Finally, a stochastic reachability problem subject to this risk metric is formulated as an LP problem to synthesize risk-aware driving policy. Such a policy can balance different types of risks, including the collision risk and the violation risk of traffic rules and norms, in terms of their timing and severity. The effectiveness and scalability of the proposed method are validated in three distinct driving scenarios using the high-fidelity Carla simulator [35], demonstrating its capability to handle complexities from traffic rules and unexpected events due to uncertainty, and its scalability across versatile scenarios.
@@ II-A Finite-State Model for Vehicles and Environments
MDP is a mathematical framework for decision-making in stochastic environments. Hence, a finite-state MDP is leveraged to capture the uncertain behavior of vehicles.
@@
We denote a finite path of $\mathcal{M}$ as $\boldsymbol{s}_{n}\!:=\!s_{0}a_{0},\ldots s_{n-1}a_{n-1}s_{n},n\!\in\!\mathbb{N}$, with $s_{i}\!\in\!S$ and $a_{i}\!\in\!A$, for all $i\!\leq\!n$. Similarly, we denote an infinite paths as $\mathbf{s}=s_{0}a_{0}s_{1}a_{1}\ldots$. The inputs of $\mathcal{M}$ are selected according to a policy defined as follows.
@@
A policy is a sequence $\pi=\pi_{0}\pi_{1}\pi_{2}\ldots$ of stochastic kernels $\pi_{n}\!:\!H\!\times\!A\!\rightarrow\![0,1]$, where $H$ denotes the collection of all the finite paths $\boldsymbol{s}_{n}$. The set of all policies is denoted by $\Pi$.
@@
Since the environmental states are uncontrollable, it is proper to use an MC model to represent the evolution of environments [36]. To sum up, an MDP model denoted by $\mathcal{M}_{v}=(S_{v},s_{v,0},A,\mathbb{T}_{v})$ is used to represent the behavior of the ego vehicle, and an MC model denoted by $\mathcal{C}_{e}=(S_{e},s_{e,0},\mathbb{T}_{e})$ to capture the behavior of the environment. Furthermore, The composition of $\mathcal{M}_{v}$ and $\mathcal{C}_{e}$ is also an MDP model, defined as ¯$\overline{\mathcal{M}}=(\bar{S},\bar{s}_{0},A,\bar{\mathbb{T}})$, where $\bar{S}\!:=\!S_{v}\!\times\!S_{e}$, $\bar{s}_{0}\!=\!(s_{v,0},s_{e,0})$, and $\bar{\mathbb{T}}\!:\!\bar{S}\!\times\!A\!\times\!\bar{S}\!\rightarrow\![0,1]$ based on $\bar{\mathbb{T}}(\bar{s},a,\bar{s}^{\prime})=\mathbb{T}_{v}(s_{v},a,s_{v}^{\prime})\times\mathbb{T}_{e}(s_{e},s_{e}^{\prime})$.
@@ II-B Specifications: Linear Temporal Logic (LTL)
Fragments: Two notable fragments of LTL are the safety and co-safety temporal logic fragments [34], which allows reasoning about the satisfaction of an infinite word using only a finite prefix. Specifically, a co-safety formula is satisfied if there exists a finite “good” prefix regardless of any infinite suffix. Conversely, the violation of a safety specification depends on the existence of a finite bad prefix. In other words, safety fragments represent “bad thing never happens”, while co-safety fragments depict “good thing eventually happens”. A syntactically co-safety LTL (scLTL) formula $\psi$ over a set of atomic propositions $\mathsf{AP}$ is recursively defined
@@
Study Case: In an intersection scenario depicted in Fig. 1, the ego vehicle is expected to eventually reach a target area labeled “$t$”, avoid collisions with the other vehicle labeled “$v$”, refrain from entering non-drivable areas labeled “$n$”, and also wait for a green light event labeled “$g$” before entering the intersection area labeled “$i$”. In this scenario, the LTL specification is $\psi=\psi_{cs}\wedge\psi_{s}$ with a co-safety formula $\psi_{cs}=\lozenge(t)$ and a safety formula $\psi_{s}=\square(\neg g\!\rightarrow\!\neg i)\wedge\square(\neg n\wedge\neg v)$, defined over $\mathsf{AP}=[t,v,g,i]$. DFAs can be autonomously translated from $\psi_{s}$ and $\psi_{cs}$ using off-the-shelf toolboxes [37].
@@ II-C Risk-Aware Control Synthesis Problem
Since traffic rules are primarily concerned with the occurrence of good or bad events rather than repetitive behaviors, we believe that the combination of safety and co-safety fragments can effectively represent nearly all traffic tasks and rules while mitigating the computational demand of control synthesis. Thus, we consider an LTL specification in the form of $\psi=\psi_{cs}\wedge\psi_{s}$ defined over an atomic proposition set $\mathsf{AP}$, composed with a safety formula $\psi_{s}$ and a co-safety formula $\psi_{cs}$. The corresponding DFAs of $\psi_{cs}$ and $\psi_{s}$ are denoted by $\mathcal{A}_{cs}$ and $\mathcal{A}_{s}$, respectively. To represent the uncertain behavior of both the environment and vehicle, we use a composed MDP $\overline{\mathcal{M}}$, as defined in Definition 1. This composed MDP combines an MDP model for an ego vehicle, $\mathcal{M}_{v}$ defined in Def. 1, and an MC model for an uncontrollable environment, $\mathcal{C}_{e}$ defined in Definition 3. The states of $\overline{\mathcal{M}}$ are labeled by $\mathcal{L}$ to an alphabet set $\Sigma$, which can be used to verify the satisfaction of $\psi$. To handle the uncertainty in realistic traffic like a human driver, we aim to develop a human-like risk metric for $\psi$ and then synthesize a stochastic policy $\pi\!\in\!\Pi$ that respects a human-like risk metric.
@@
Risk metrics are often defined based on simple reach-avoid problems [38]. Therefore, we will start by rewriting the satisfaction problem for the given specifications as a reach-avoid problems over product MDP.
@@ III-B Limitations of Event Probability
However, using satisfaction and violation probability defined in (2) and (3) to evaluate safety in realistic traffic might lead to unnatural behavior. The event probability in (2) and (3) means that each possible trace generated by $\pi$ that drives $\mathcal{P}$ to enter into final states in finite time gets weight 1 and all else gets weighted 0, which is displayed in Fig. 2. Such an approach embraces two points against the awareness way of human drivers: (1) It does not account for the timing of violations, which may cause the vehicle to react too early to dangers that happen in the distant future; (2) It fails to consider varying costs associated with various types of violations, preventing the vehicle from appropriately balancing the seriousness of various dangers. Thus, a risk metric reflecting human-like risk awareness is necessary.
@@
III-C Using Discounting to Design Human-like Metrics
When rewards for decisions occur at different points in the future, discounted rewards are commonly used in reinforcement learning and economics. The discounted version of satisfaction probability, $\mathbb{P}_{\pi}(\boldsymbol{\omega}\models\psi_{cs})$, can be defined as,
@@ III-C Using Discounting to Design Human-like Metrics
where $\gamma<1$ is a discount factor. This discounting represents the preference of people for obtaining rewards sooner rather than later. The rationality of the discounting operation has also been supported by psychology studies [39]. When it comes to violation risks of safety specification, we propose a similar metric. In addition, varying costs of events are considered by a cost mapping function, denoted as $c(\cdot)\!:\!Z\!\rightarrow\!\mathbb{R}$, that assigns cost values to different product states. Specifically, $c(\hat{z})=0$ if $\hat{z}\!\notin\!D$, while $c(\hat{z})>0$ if $\hat{z}\!\in\!D$. Values of $c(\hat{z}\!\in\!D)$ is associated with ethical costs in real life. Furthermore, we propose the following metric to evaluate the risk of a policy $\pi$,
@@
The proposed metric can conquer the limitations of event probabilities. As depicted in Fig. 3, discounting reflects how a driver’s awareness gradually decreases over time, causing the perceived risk for the same event to also diminish over time. When an event is significantly severe, like the one indicated by the red arrow, the risk metric in (5) will attach importance to this event more than less severe events. Therefore, this metric is able to consider the timing of events and also balance different severity of risks. In addition, this risk metric can be seen as a formaization of the concept of the driver’s risk field (DRF) model [24, 28], which is widely used in the transportation field.
@@
The Driver’s Risk Field (DRF) model is designed empirically to evaluate the uncertainty distribution from a driver’s view. The risk field $\alpha(\cdot)\!:\!\mathbb{R}^{2}\!\rightarrow\!\mathbb{R}$ maps horizontal coordinates to scalar values representing the driver’s attention. To model human-like awareness, the DRF model operates on the intuition that the importance of future events diminishes while the corresponding uncertainty increases along the predictive horizon. The detailed formulation of DRF can be found in [24]. Subsequently, perceived risk can be evaluated via the expectation of the event cost, i.e., the product of the risk field and the cost map $c(\cdot)\!:\!\mathbb{R}^{2}\!\rightarrow\!\mathbb{R}$ that represents the cost distribution over horizontal coordinates. Given a griding discretization, the risk metric is defined as:
@@
where $X$ and $Y$ are finite sets consisting of discretized coordinates along with $x$ and $y$-axis, respectively.
Inspired by the risk metric of DRF in Eq. (6) , the risk metric in Eq. (5) can be rewritten as follows:
@@
In this formulation, $\beta_{\bar{\pi}}(z)$ is a discounted occupation measure. An occupation measure describes the expected amount of times that a state has been visited before reaching the final set given a Markove and stationary policy $\bar{\pi}(z,a)$, formulated as $\bar{\beta}_{\bar{\pi}}(z):=\mathbb{E}_{\bar{\pi}}\left[\sum_{t=0}^{\infty}\mathbf{1}\left(z_{t}=z\right)\right]$. Occupation measures can be used to formulate stochastic shortest path problems [31]. By comparing risk metrics in (7) and (6), it is clear that the discounted occupation measure plays a similar role as the DRF model. In addition, the discounted occupation measure embraces similar awareness for events as the DRF model. Thus, we conjecture that the discounted occupation measure is equivalent to the human-like risk awareness model, DRF. Thus, we can integrate DRF into a control synthesis problem for LTL specification using discounted occupation measures.
IV Synthesis of human-like risk-aware control
In this section, we formulate a reasonable control synthesis problem to trade off the rewards metric for co-safety specification and the risk metric for safety specification. A reasonable approach is to maximize the rewards metric and also keep the risk metric in an acceptable range, which is formulated as follows,

### v2

# Risk-Aware Autonomous Driving with Linear Temporal Logic Specifications

@@ Abstract
Humans naturally balance the risks of different concerns while driving, including traffic rule violations, minor accidents, and fatalities. However, achieving the same behavior in autonomous systems remains an open problem. This paper extends a risk metric that has been verified in human-like driving studies to encompass more complex driving scenarios specified by linear temporal logic (LTL) that go beyond just collision risks. This extension incorporates the timing and severity of events into LTL specifications, thereby reflecting a human-like risk awareness. Without sacrificing expressivity for traffic rules, we adopt LTL specifications composed of safety and co-safety formulas, allowing the control synthesis problem to be reformulated as a reachability problem. By leveraging occupation measures, we formulate a linear programming (LP) problem for this LTL-based risk metric. Consequently, the synthesized policy balances different types of risks, including not only collision risks but also traffic rule violations. The effectiveness of the proposed approach is validated by three typical traffic scenarios in the Carla simulator.
@@ I Introduction
Developing motion planning frameworks for autonomous vehicles remains a significant challenge due to the inherent complexity and uncertainty of real-world traffic [1, 2]. The complexity arises primarily from the diverse concerns in practical driving tasks, including safety [3], traffic rules [4], and social norms [5]. The uncertainty mainly originates from the dynamic environment with unpredictable traffic participants [6] and unexpected events [7]. Taking an unprotected turn scenario in Fig. 1 as an example, the ego vehicle must make a left turn while avoiding collisions with the opponent car, obey traffic rules such as reacting to traffic lights, and follow driving norms like yielding to oncoming vehicles. Environmental uncertainty imposes the ego vehicle at the potential risk of violating the above specifications. Therefore, a risk-aware autonomous driving scheme is essential to make decisions that balance diverse risks associated with safety, traffic rules, and social norms [8].
@@
Temporal logic provides a formal and expressive framework for describing high-level tasks that incorporate spatial and temporal constraints and synthesize provably correct controllers [9, 10, 11], which makes it well suited for encoding traffic tasks with rules and norms [12] and improving trustworthiness in motion planning tasks [13, 14]. While temporal logic-based planning offers significant advantages, existing frameworks remain limited and impractical for real-world traffic deployment. Due to uncertainty and dynamic interaction in real traffic, achieving complete safety and full satisfaction of temporal logic specifications is unrealistic, making it essential to develop a risk-aware framework. Existing temporal logic planning methods quantify risk as the probability that a temporal logic specification fails and focus on maximizing the satisfaction probability of these specifications [15, 16, 17, 18, 19]. However, such risk metrics are not suitable for realistic traffic where different types of risky events may be involved [20]. In practical traffic with diverse risks, human drivers intuitively prioritize near-future and more severe events over those in the distant future or with less severity [21]. Unfortunately, the satisfaction probability of temporal logic specifications fails to differentiate the severity and timing of events [22], which results in illogical and even unsafe driving behavior.
The driver’s risk field model [23, 24], proposed by the transportation community, has been shown to align with human-like risk awareness by discounting collision risks over the prediction horizon. This mode enables vehicles to track references safely by balancing risks between multiple obstacles. However, this model is limited to collision avoidance in tracking tasks, making it less effective for handling realistic traffic scenarios with complex rules. To address this, we extend the risk field model by integrating temporal logic specifications. This paper selects to use LTL [9], a classic type of temporal logic widely used in motion planning. Specifically, we extend upon the tracking task with more complex driving goals expressed by co-safe specifications and quantify risk specified as breaching safety specifications. By incorporating the inherent ability to balance timing and severity from the risk field model, this extension overcomes the limitation of temporal logic mentioned above. Safety or co-safety specifications [25] are special fragments of LTL that can transform the control synthesis problem into a classical reachability problem. Therefore, this approach retains the ability to simplify the LTL planning into a reachability problem while preserving the necessary expressivity for traffic tasks. To formulate a tractable control synthesis problem, we rewrite the risk metric using the discounted version of occupation measures, a key concept for solving a stochastic reachability problem for a Markov Decision Process (MDP) [26]. Referring to the multi-objective control synthesis formulation for LTL in [27], we formulate a linear programming (LP) problem over product automaton to synthesize risk-aware policies using occupation measures as decision variables.
The effectiveness and scalability of the proposed method are validated in three distinct driving scenarios using the high-fidelity Carla simulator [28], demonstrating its capability to handle complexities from traffic rules and unexpected events due to uncertainty, and its scalability across versatile scenarios.
@@ II-A Finite-State Model for Vehicles and Environments
MDP is a typical mathematical framework for decision-making in stochastic environments, and its variants, such as the partially observable MDP, can characterize more complex traffic scenarios [29]. As the first step before considering more complex models, this paper begins by using a finite-state MDP to capture the uncertain behavior of vehicles.
@@
We denote a finite path of $\mathcal{M}$ as $\boldsymbol{s}_{n}\!:=\!s_{0}a_{0}\ldots a_{n-1}s_{n}$, $n\!\in\!\mathbb{N}$, where $s_{i}\!\in\!S$ and $a_{i}\!\in\!A$ for all $i\!\leq\!n$. Similarly, we denote an infinite paths as $\mathbf{s}\!=\!s_{0}a_{0}s_{1}a_{1}\ldots$. The inputs of $\mathcal{M}$ are selected according to a policy defined as follows.
@@
A policy is a sequence $\pi\!:=\!\pi_{0}\pi_{1}\pi_{2}\ldots$ of stochastic kernels $\pi_{n}\!:\!H\!\times\!A\!\rightarrow\![0,1]$, where $H$ denotes the collection of all the finite paths $\boldsymbol{s}_{n}$. The set of all policies is denoted by $\Pi$.
@@
Since environmental states are uncontrollable, an MC is used to represent the evolution of environments [30]. In summary, an MDP, denoted by $\mathcal{M}_{v}\!=\!(S_{v},s_{v}^{0},A,\mathbb{T}_{v})$, is used to represent the behavior of the ego vehicle, while an MC model, denoted by $\mathcal{C}_{e}=(S_{e},s_{e}^{0},\mathbb{T}_{e})$, captures the behavior of the environment. Furthermore, the composition of $\mathcal{M}_{v}$ and $\mathcal{C}_{e}$ forms a new MDP model, defined as ¯$\overline{\mathcal{M}}=(\bar{S},\bar{s}_{0},A,\bar{\mathbb{T}})$, where $\bar{S}\!:=\!S_{v}\!\times\!S_{e}$, $\bar{s}_{0}\!=\!(s_{v}^{0},s_{e}^{0})$, and $\bar{\mathbb{T}}\!:\!\bar{S}\!\times\!A\!\times\!\bar{S}\!\rightarrow\![0,1]$ based on $\bar{\mathbb{T}}(\bar{s},a,\bar{s}^{\prime})=\mathbb{T}_{v}(s_{v},a,s_{v}^{\prime})\times\mathbb{T}_{e}(s_{e},s_{e}^{\prime})$.
@@ II-B Specifications: Linear Temporal Logic (LTL)
Fragments: Two notable fragments of LTL are the safety and co-safety specifications [25], which allows reasoning about the satisfaction of an infinite word using only a finite prefix. Specifically, a co-safety formula is satisfied if there exists a finite “good” prefix regardless of any infinite suffix. Conversely, the violation of a safety specification depends on the existence of a finite bad prefix. In other words, safety fragments represent “bad thing never happens”, while co-safety fragments depict “good thing eventually happens”. A syntactically co-safety LTL (scLTL) formula $\psi$ over a set of atomic propositions $\mathsf{AP}$ is recursively defined as,
@@
Study Case: In an intersection scenario depicted in Fig. 1, the ego vehicle is expected to eventually reach a target area labeled “$t$”, avoid collisions with the other vehicle labeled “$v$”, refrain from entering non-drivable areas labeled “$n$”, and also wait for a green light event labeled “$g$” before entering the intersection area labeled “$i$”. In this scenario, the LTL specification is $\psi=\psi_{cs}\wedge\psi_{s}$ with a co-safety formula $\psi_{cs}=\lozenge(t)$ and a safety formula $\psi_{s}=\square(\neg g\!\rightarrow\!\neg i)\wedge\square(\neg n\wedge\neg v)$, defined over $\mathsf{AP}\!=\!\{t,v,g,i\}$. DFAs can be autonomously translated from $\psi_{s}$ and $\psi_{cs}$ using off-the-shelf toolboxes [31].
@@ II-C Risk-Aware Control Synthesis Problem
Since traffic rules are primarily concerned with the occurrence of good or bad events rather than repetitive behaviors, we use the combination of safety and co-safety fragments to effectively represent nearly all traffic tasks and rules while mitigating the computational demand of control synthesis. Thus, we consider an LTL specification in the form of $\psi=\psi_{cs}\wedge\psi_{s}$ defined over an atomic proposition set $\mathsf{AP}$, composed with a safety formula $\psi_{s}$ and a co-safety formula $\psi_{cs}$. The corresponding DFAs of $\psi_{cs}$ and $\psi_{s}$ are denoted by $\mathcal{A}_{cs}$ and $\mathcal{A}_{s}$, respectively. To represent the uncertain behavior of both the environment and vehicle, we use a composed MDP $\overline{\mathcal{M}}$, as defined in Definition 1. This composed MDP combines an MDP model for an ego vehicle, $\mathcal{M}_{v}$ defined in Def. 1, and an MC model for an uncontrollable environment, $\mathcal{C}_{e}$ defined in Definition 3. The states of $\overline{\mathcal{M}}$ are labeled by $\mathcal{L}$ to an alphabet set $\Sigma$, which can be used to verify the satisfaction of $\psi$. To handle the uncertainty in realistic traffic like a human driver, we aim to develop a human-like risk metric for $\psi$ and then synthesize a stochastic policy $\pi\!\in\!\Pi$ that respects a human-like risk metric.
@@
Risk metrics are often defined based on simple reach-avoid problems [32]. Therefore, we will start by rewriting the satisfaction problem for the given specifications $\psi=\psi_{cs}\wedge\psi_{s}$ as a reach-avoid problem over product MDP.
@@ III-B Limitations of Event Probability
Using satisfaction and violation probability defined in (2) and (3) to evaluate safety in realistic traffic might lead to unnatural behavior. The event probability in (2) and (3) means that each possible trace generated by $\pi$ that drives $\mathcal{P}$ to enter into final states in finite time gets weight 1 and all else gets weighted 0, which is displayed in Fig. 2. Such an approach embraces two points against the awareness way of human drivers: (1) It does not account for the timing of violations, which may cause the vehicle to react too early to dangers that happen in the distant future; (2) It fails to consider varying costs associated with various types of violations, preventing the vehicle from appropriately balancing the seriousness of various dangers. Thus, a risk metric reflecting human-like risk awareness is necessary.
@@
III-C Human-like Metric Design
Discounted rewards are commonly used in reinforcement learning and economics to account for the timing of events when decisions lead to rewards at different moments in the future. The satisfaction probability, $\mathbb{P}_{\pi}(\boldsymbol{\omega}\models\psi_{cs})$, can be modified by introducing discounting as follows,
@@ III-C Human-like Metric Design
where $\gamma<1$ is a discount factor. This discounting represents the preference of people for obtaining rewards sooner rather than later. The rationality of the discounting operation has also been supported by psychology studies [33]. When it comes to violation risks of safety specification, we propose a similar metric. In addition, varying severity of events are considered by a cost mapping function, denoted as $c(\cdot)\!:\!Z\!\rightarrow\!\mathbb{R}$, that assigns cost values to different product states. Specifically, $c(\hat{z})=0$ if $\hat{z}\!\notin\!D$, while $c(\hat{z})>0$ if $\hat{z}\!\in\!D$. Furthermore, we propose the following metric to evaluate the risk of a policy $\pi$,
@@
The proposed metric addresses the limitations of event probabilities. As depicted in Fig. 3, discounting reflects how a driver’s awareness gradually decreases over time, causing the perceived risk for the same event to decrease accordingly. When an event is significantly severe, like the one indicated by the red arrow, the risk metric in (5) places greater importance on this event than on less severe ones. Therefore, this metric can account for both the timing and severity of risky events. Defined over product state space, this metric can be seen as an extension of the driver’s risk field (DRF) model [23, 24] to LTL specifications.
@@
Driver’s Risk Field (DRF) model is designed empirically to evaluate the uncertainty distribution from a driver’s view. The risk field $\alpha(\cdot)\!:\!\mathbb{R}^{2}\!\rightarrow\!\mathbb{R}$ maps horizontal coordinates to scalar values representing the driver’s attention. To model human-like awareness, DRF is designed based on the intuition that the importance of future events diminishes while the corresponding uncertainty increases along the predictive horizon. A detailed formulation of DRF can be found in [23]. Perceived risk is evaluated by calculating the expected event cost, which is the product of the risk field and the cost map. With a grid discretization, the DRF risk metric is defined as:
@@
where $c(\cdot)\!:\!\mathbb{R}^{2}\!\rightarrow\!\mathbb{R}$ represents the cost distribution over horizontal coordinates, and $X$ and $Y$ are finite sets of discretized coordinates along with $x$ and $y$-axis.
Inspired by DRF risk metric in Eq. (6) , the risk metric in Eq. (5) is rewritten as follows:
@@
Referring to the standard occupation measure formulated as $\mathbb{E}_{\bar{\pi}}\textstyle\left[\sum_{t}^{\infty}\mathbf{1}(z_{t}=z)\right]$, $\beta_{\bar{\pi}}(z)$ can be regarded as a discounted version of the occupation measure. An occupation measure describes the expected number of times that a state has been visited before reaching the final set, given a Markove and stationary policy $\bar{\pi}(z,a)$. Occupation measures can be used to formulate stochastic shortest path problems [34]. By comparing risk metrics in (7) and (6), it is evident that the discounted occupation measure plays a similar role as the DRF model. In addition, the discounted occupation measure embraces similar awareness for events as the DRF model. Thus, we conjecture that the proposed risk metric embodies a similar human-like risk awareness as DRF.
IV Risk-aware control
In this section, we formulate a first risk-aware control synthesis problem with the proposed metrics. A reasonable approach is to maximize the rewards metric of co-safety specification and also keep the risk metric of safety specification in an acceptable range, which is formulated as follows,